# Anomalous dynamical Casimir effect in an expanding ring

### v2

@@ Abstract
Supplemental Materials: Analog dynamical curved spacetimes in Bose-Einstein condensates
@@ I Curved spacetimes in Bose-Einstein condensates
In this appendix, following the review of Barcelo, Liberati, and Visser Barcelo et al. (2001) we analyze the possibility to generate analog curved spacetimes in Bose-Einstein condensates.
@@ I.1 From Bose-Einstein condensates to hydrodynamics-like equations
where $\psi(\vec{x},t)=\langle\hat{\Psi}(\vec{x},t)\rangle$ is referred to as the wave function of the condensate Bogoliubov (1947), while $\hat{\phi}$ denotes the excitation on top of the condensate.
@@
They are identical to Euler equations for a barotropic inviscid irrotational fluid Unruh (1981) of velocity $\vec{v}$ with $\vec{v}\equiv\frac{1}{m}\vec{\nabla}\theta_{0}$ in the presence of a quantum potential term $V_{quant}=-\frac{\hbar^{2}}{2m}\frac{\vec{\nabla}^{2}\sqrt{n_{0}}}{\sqrt{n_{0}}}$.
@@ I.2 Sound propagation and metric tensor
Starting from equation (S4), let us show that in a way similar to the classical hydrodynamic, studied by Unruh Unruh (1981), the weak fluctuation of $\psi$ around a background $\psi_{0}$ can be described by a curved spacetime Klein-Gordon equation. To do so, let us develop $\psi$ around a background $\psi_{0}=\sqrt{n_{0}}\exp\left(i\theta_{0}/\hbar\right)$ such as
@@
In 1+1 dimensions, due to the conformal symmetry, every metric verifies $\det\left(\sqrt{-\det(g_{\rho\sigma})}g^{\mu\nu}\right)=1$. Therefore, whenever $\det\left(f^{\mu\nu}\right)$ is not a constant (S10) does not describe the dynamics of a massless Boson in curved spacetime [^1] If it is a constant different from one can simply factor it out of equation (S10).. Starting from a metric $f^{\mu\nu}$ with $\det\left(f^{\mu\nu}\right)=Cst$, we identify a family of metrics $g^{\mu\nu}$ of the form $g_{\mu\nu}=e^{\sigma(x)}f_{\mu}\nu$ that satisfy
@@ II Anomalous stress-energy tensor in time-dependent metrics
In 1977, Christensen and Fulling Christensen and Fulling (1977) computed the stationary momentum energy tensor for both a 3+1 and a 1+1 dimensional Schwarzschild black hole. In a previous article Bermond et al. (2022), we generalized their argument to determine the stationary momentum energy tensor of a chiral field in any background 1+1 dimensional static metric. Here, we reconsider these arguments in the case of a 1+1 dimensional dynamical metric of the form
@@ II.1 Gravitational anomalies
Focusing on a symmetric momentum energy tensor, the anomalies, encoding the violation of classical symmetries by quantum fluctuations in the presence of a background metric, can be written Capper and Duff (1974); Capper et al. (1974); Bertlmann (2000) as
@@
with $c$ the central charges and $\chi$ the chiralities. In other words, this corresponds to $\mathcal{C}_{g}=c_{+}-c_{-}$ and $\mathcal{C}_{w}=c_{+}+c_{-}$. A simple proof of the formula (S19a,S19b) is derived by Bertlmann and Kohlprath in Bertlmann and Kohlprath (2001) for free chiral fermions with $\mathcal{C}_{w}=1$ and $\mathcal{C}_{g}=\pm 1$.
@@ II.3 Momentum energy tensor
with $\left(T^{\mu}_{\phantom{0}\nu}\right)_{0}$, the classical solution (verifying (II.3) with $\mathcal{R}=0$)
@@ III Anomalous dynamical Casimir effect for a Bose-Einstein condensate time-dependent velocity
In the main text, we considered a situation in which one induces an effective curvature of spacetime in a Bose-Einstein condensate by modulating the trapping shape. Another experimentally relevant strategy to induce a non-zero spacetime curvature consists in modulating over time the sound velocity of the condensate. Such a variation can in principle be implemented by resorting to Feshbach resonances Vogels (1999); Inouye et al. (1998) and leads to a dynamical metric of the form
@@
The finite size instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)=\frac{\pi\hbar c_{s}(t)}{6L^{2}}$ is therefore modified by the presence of an external acceleration $\partial_{t}c_{s}$. In particular, the energy density is reminiscent of the result obtained by W.Unruh in a different contextUnruh (1976). Indeed, defining $T_{U}$ as the Unruh temperature such as
@@
wich corresponds to a quarter of Unruh thermal energy, $\varepsilon_{U}=\mathcal{C}_{w}\gamma T_{U}^{2}/2$ Fulling (1973); Davies (1975); Unruh (1976). Therefore, the total energy density is the sum of the instantaneous Casimir energy and the instantaneous Unruh energy density.
While Eq. (III) gives us access to the anomalous correction for any velocity ramp profile, in this section, for simplicity, we will consider a ramp of the sound velocity of the form
@@
The parameters of Fig. S1 are the same that those used for Fig. 2, motivated by the realization of a Bose-Einstein condensate in ${}^{23}$a by S. Eckel et al. described in Eckel et al. (2018). For a slow ramp of the condensate scattering lengths, corresponding to a ramp over a time $\tau=15ms$, $\varepsilon_{\bar{\mathcal{R}}}$ is negligible, such that there is a single anomalous energy scale. As a consequence, the corrected energy density is approximately equal to its classical value, the instantaneous Casimir energy $\varepsilon_{\mathcal{C}}(t)$, while the pressure slightly deviates from this classical value, with a deviation quantified by $\varepsilon_{\mathcal{R}}$.

### v3

@@ Acknowledgements.
SUPPLEMENTAL MATERIAL
Anomalous dynamical Casimir effect in an expanding ring
@@ I Curved spacetimes in Bose-Einstein condensates
In this appendix, following the review of Barcelo, Liberati, and Visser barcelo2001analogue we analyze the possibility to generate analog curved spacetimes in Bose-Einstein condensates.
@@ I.1 From Bose-Einstein condensates to hydrodynamics-like equations
where $\psi(\vec{x},t)=\langle\hat{\Psi}(\vec{x},t)\rangle$ is referred to as the wave function of the condensate bogoliubov1947theory, while $\hat{\phi}$ denotes the excitation on top of the condensate.
@@
They are identical to Euler equations for a barotropic inviscid irrotational fluid unruh1981experimental of velocity $\vec{v}$ with $\vec{v}\equiv\frac{1}{m}\vec{\nabla}\theta_{0}$ in the presence of a quantum potential term $V_{quant}=-\frac{\hbar^{2}}{2m}\frac{\vec{\nabla}^{2}\sqrt{n_{0}}}{\sqrt{n_{0}}}$.
@@ I.2 Sound propagation and metric tensor
Starting from equation (S4), let us show that in a way similar to the classical hydrodynamic, studied by Unruh unruh1981experimental, the weak fluctuation of $\psi$ around a background $\psi_{0}$ can be described by a curved spacetime Klein-Gordon equation. To do so, let us develop $\psi$ around a background $\psi_{0}=\sqrt{n_{0}}\exp\left(i\theta_{0}/\hbar\right)$ such as
@@
In 1+1 dimensions, due to the conformal symmetry, every metric verifies $\det\left(\sqrt{-\det(g_{\rho\sigma})}g^{\mu\nu}\right)=1$. Therefore, whenever $\det\left(f^{\mu\nu}\right)$ is not a constant (S10) does not describe the dynamics of a massless Boson in curved spacetime. Starting from a metric $f^{\mu\nu}$ with $\det\left(f^{\mu\nu}\right)=Cst$, we identify a family of metrics $g^{\mu\nu}$ of the form $g_{\mu\nu}=e^{\sigma(x)}f_{\mu}\nu$ that satisfy
@@ II Anomalous stress-energy tensor in time-dependent metrics
In 1977, Christensen and Fulling christensen1977trace computed the stationary momentum energy tensor for both a 3+1 and a 1+1 dimensional Schwarzschild black hole. In a previous article bermond2022anomalous, we generalized their argument to determine the stationary momentum energy tensor of a chiral field in any background 1+1 dimensional static metric. Here, we reconsider these arguments in the case of a 1+1 dimensional dynamical metric of the form
@@ II.1 Gravitational anomalies
Focusing on a symmetric momentum energy tensor, the anomalies, encoding the violation of classical symmetries by quantum fluctuations in the presence of a background metric, can be written capper1974one; capper1974photon; bertlmann2000anomalies as
@@
with $c$ the central charges and $\chi$ the chiralities. In other words, this corresponds to $\mathcal{C}_{g}=c_{+}-c_{-}$ and $\mathcal{C}_{w}=c_{+}+c_{-}$. A simple proof of the formula (S19a,S19b) is derived by Bertlmann and Kohlprath in bertlmann2001two for free chiral fermions with $\mathcal{C}_{w}=1$ and $\mathcal{C}_{g}=\pm 1$.
@@ II.3 Momentum energy tensor
with $\left(T^{\mu}_{\phantom{0}\nu}\right)_{0}$, the classical solution (verifying (S33) with $\mathcal{R}=0$)
@@ III Anomalous dynamical Casimir effect for a Bose-Einstein condensate time-dependent velocity
In the main text, we considered a situation in which one induces an effective curvature of spacetime in a Bose-Einstein condensate by modulating the trapping shape. Another experimentally relevant strategy to induce a non-zero spacetime curvature consists in modulating over time the sound velocity of the condensate. Such a variation can in principle be implemented by resorting to Feshbach resonances vogels1999feshbach; inouye1998observation and leads to a dynamical metric of the form
@@
The finite size instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)=\frac{\pi\hbar c_{s}(t)}{6L^{2}}$ is therefore modified by the presence of an external acceleration $\partial_{t}c_{s}$. In particular, the energy density is reminiscent of the result obtained by W.Unruh in a different contextunruh1976notes. Indeed, defining $T_{U}$ as the Unruh temperature such as
@@
wich corresponds to a quarter of Unruh thermal energy, $\varepsilon_{U}=\mathcal{C}_{w}\gamma T_{U}^{2}/2$ fulling1973nonuniqueness; davies1975scalar; unruh1976notes. Therefore, the total energy density is the sum of the instantaneous Casimir energy and the instantaneous Unruh energy density.
While Eq. (S42) gives us access to the anomalous correction for any velocity ramp profile, in this section, for simplicity, we will consider a ramp of the sound velocity of the form
@@
The parameters of Fig. S1 are the same that those used for Fig. 2 of the main text, motivated by the realization of a Bose-Einstein condensate in 23Na by S. Eckel et al. described in eckel2018rapidly. For a slow ramp of the condensate scattering lengths, corresponding to a ramp over a time $\tau=15ms$, $\varepsilon_{\bar{\mathcal{R}}}$ is negligible, such that there is a single anomalous energy scale. As a consequence, the corrected energy density is approximately equal to its classical value, the instantaneous Casimir energy $\varepsilon_{\mathcal{C}}(t)$, while the pressure slightly deviates from this classical value, with a deviation quantified by $\varepsilon_{\mathcal{R}}$.